### v1

# Genetic Data Governance in Crisis: Policy Recommendations for Safeguarding Privacy and Preventing Discrimination

Abstract.
Genetic data collection has become ubiquitous today. The ability to meaningfully interpret genetic data has motivated its widespread use across forensics, clinical practice, and research, providing crucial insights into human health and ancestry while driving important public health initiatives. Easy access to genetic testing has fueled a rapid expansion of direct-to-consumer offerings, many of which are recreational in nature. However, the growth of genetic datasets and their applications has created significant privacy and discrimination risks, particularly as our understanding of the scientific basis for genetic traits continues to evolve. In this paper, we organize the uses of genetic data along four distinct ‘pillars’: clinical practice, research, forensic and government use, and recreational use. Using our scientific understanding of genetics, genetic inference methods and their associated risks, and existing regulatory mechanisms, we build a risk assessment framework that identifies key values that any governance system must preserve. We then analyze case studies from each of the pillars using this framework to assess how well existing legal and regulatory frameworks preserve desired values. Our investigation reveals critical gaps in existing regulatory frameworks and identifies specific threats to privacy and personal liberties, particularly through genetic discrimination. To address these challenges, we call for and propose comprehensive regulatory reforms including: (1) updating the legal definition of genetic data to protect against modern technological capabilities, (2) expanding the Genetic Information Nondiscrimination Act (GINA) to cover currently unprotected domains, and (3) establishing a unified regulatory framework under a single governing body to oversee all applications of genetic data.
We conclude with three open questions about genetic data: the challenges posed by its relational nature, including consent for relatives and minors; the complexities of international data transfer; and its potential integration into large language models.
Data Governance, Genetic Privacy, Regulatory Frameworks, Genetic Discrimination, Policy, Data Protection, Genetic Inference
$c$s: Applied computing Computational genomicsccs: Applied computing Health informaticsccs: Applied computing Lawccs: Applied computing Consumer health
1. Introduction
The scale of genetic data generation has increased at a staggering rate since 2001 (Bick et al., 2024; Bycroft et al., 2018; Loos, 2020). A primary driver of this is decreased DNA sequencing costs (Wetterstrand, 2023), which have enabled the curation of large genetic datasets for personal and public health, ancestry inference, relative identification and advanced forensics. For example, there are whole genome sequences combined with medical records from approximately 400,000 US residents in the National Institutes of Health (NIH) All of Us Research Study, with plans to increase sampling to include at least 1 million US residents (National Institutes of Health, 2024). Direct-to-consumer (DTC) genetic testing companies have much more data: 23andMe has 14 million customer sequences (23a, 2023a) and Ancestry has 25 million (Ancestry, 2025). The increase in genetic data and improvements in technologies has resulted in the following aspirational “bold prediction” for 2030 by the NIH: “a person’s complete genome sequence along with informative annotations can be securely and readily accessible on their smartphone” (Green et al., 2020). From a scientific perspective, the collection of genetic data at this scale has been critically valuable to understanding health, ancestry, and driving public health initiatives.
The propensity for genetic discrimination grows with poor protections for privacy. Since 2020, the U.S. Department of Homeland Security has accelerated its collection of genetic data, growing the CODIS database by more than 1.5 million people, the majority of whom are people of color (Glaberson et al., 2024). 23andMe—which boasts more than 14 million customers (23a, 2023a)—suffered a data breach after a hacker accessed 14,000 customer profiles [^1] While genetic data was not accessed, ancestry information and relative names were(23a, 2023b), primarily targeting individuals of Chinese and Ashkenazi Jewish descent (Carballo et al., 2024). These pressing, real-world discrimination and privacy concerns are exacerbated by the sordid history of eugenics and forced sterilization in U.S. states in the 20th century (a model later adopted by Nazi Germany (Spiegel, 2019)), and more recent practices of genetic determinism (Epstein, 2003), racial pseudoscience (Lala and Feldman, 2024; Duello et al., 2021; Harmon, 2018; Panofsky et al., 2021), race-motivated violence (Carlson et al., 2022), privacy and consent violations (Strand, 2016; Sterling, 2011; Jillson, 2024), and broad scientific distrust (Panacer, 2023; Kaye, 2012; Saulsberry and Terry, 2013). These are serious concerns that must be addressed today given the number of recent proposals for genetic data applications in public infrastructure, (e.g., to assess intelligence proxies, like IQ (R and S, 2018); insurance premiums (Karlsson Linnér and Koellinger, 2022); or educational outcomes (Harden et al., 2020)). The only federal-level legislation currently in effect is the Genetic Information Non-Discrimination Act (GINA) of 2008. GINA prohibits the use of genetic information in employment decisions and discrimination based on genetic information in health insurance coverage but does not prohibit genetic discrimination for long-term care, disability, or life insurance.
Our Work.
In this paper, we make three contributions towards the construction of a robust genetic data governance system. First, we expand upon Wan et al.’s (Wan et al., 2022) Four Pillars of genetic data collectors and make the argument that inconsistent and leaky regulation across these pillars introduces opportunity for genetic discrimination and privacy violations. Second, we outline a risk assessment framework for modern-day genetic data governance and outline key values it should preserve. Third, we make three concrete recommendations to: (1) legally redefine genetic data; (2) expand GINA; and (3) create a regulatory framework for genetic data collectors.
Our paper is structured as follows: In Section 2, we provide an overview of genetic data and inference, with a focus on the nuances of genetic trait prediction. In Section 3, we introduce the extended Four Pillars and describe the existing legislative and regulatory structure governing genetic data use. In Section 4, we outline our risk assessment framework and apply it to four case studies in Section 5. We give our three policy recommendations in Section 6 and conclude, in Section 7, with three open challenges that are beyond the scope of this paper.
2. Background: How meaning is derived from genetic data
In this section, we provide a brief overview of genetic data and genetic inference. An individual’s genome is over 3 billion DNA bases [^2] Units of DNA long and derived equally from both genetic parents. Genetic variants—units of DNA that differ between individuals—are scattered throughout the genome and contain information about genetically influenced traits (from physical traits to behavioral traits). To predict traits or disease risk, geneticists analyze an individual’s unique set of genetic variants. However, this remains a challenging task: while diseases like cystic fibrosis or Huntington’s disease are determined by a single genetic variant, most traits are influenced by many variants across the genome as well as non-genetic factors (complex traits).
2.1. Your genetic data is shaped by your ancestors and shared with your relatives.
The inheritance of DNA—where each parent passes down a random half of their genome to their offspring—results in relatives sharing portions of their DNA, with closer relatives sharing a larger proportion. This shared genetic material means that decisions made about any individual’s genome inherently extend to their genetic relatives. This relational nature of genetic data—the genetic dragnet—has been noted to significantly complicate discussions about privacy (Costello, 2022). A powerful example of the genetic dragnet came in 2018, when law enforcement leveraged it to identify the Golden State Killer. Authorities used DNA left at a crime scene to identify distant relatives through DNA matching on a public website, GEDMatch. By combining these matches with publicly available genealogical records, they identified the suspect despite the closest identifiable relatives sharing only a great-great-great-great-grandfather (Kaiser, 2018; Zabel, 2019). At the same time, an individual’s genome is unique to them: even the genomes of identical twins contain differences (Ormond et al., 2024). While a single genome in isolation may not always be immediately identifiable the relational nature of genetic data can enable re-identification. Researchers have argued that genomic data should be treated as ”always identifiable in principle” due to this inherent interconnectedness (Shabani and Marelli, 2019; Bonomi et al., 2020).
2.2. Geneticists use statistical models for trait prediction.
Inference using genetic data requires some knowledge of a variant’s effect on a trait or disease. This can sometimes be tested in animals by controlled experiments that manipulate the animal’s genome. However, this approach is infeasible and unethical to pursue in humans, and so researchers instead rely on large genetic datasets to ask ”Do individuals who have the variant tend to have the disease (or trait of interest) compared to those who do not have the variant?” To answer this question, researchers typically conduct genome-wide association studies (GWAS), which identify genetic variants that are associated with a trait and quantify the effect of each variant on the trait. The learned effect sizes of variants (also called weights) are directly interpretable [^3] For example, in a GWAS for height, an effect size of 0.01 would mean that the variant increases height by an average of 0.01 cm. and can be used for trait prediction. The polygenic score, considered the gold standard for trait prediction, is calculated as a weighted sum of genetic variants, with each variant scaled by its corresponding GWAS effect size. See Section 7 for a discussion of frontier models for genetic data (Fritzsche et al., 2023).
2.3. Insight from genetic data is a function of the cohort.
The reference genetic dataset used for a GWAS is analogous to the training dataset in machine learning. Reference datasets contain genetic data and their “ground truth” annotations of trait values, whether it be ancestry, disease status, height, etc. Non-genetic factors of a cohort also influence a GWAS. Confounders like environmental exposures and social determinants of health can introduce bias in a genetic study. For example, the average participant in the UK Biobank  [^4] The UK Biobank is a leading resource which contains genetic and medical data from 500,000 UK residents aged 40-69 at the time of recruitment. tends to be more educated and healthier than the average Briton (Fry et al., 2017). The choice of the cohort—that is who is in the dataset and how their data is labeled555‘True” ancestry has been arbitrarily (and implicitly) set by the field to correspond to: where your ancestors lived 600 years ago (Coop, 2023). Individuals in the same ancestry group are distantly related in the same timescale (i.e., likely many ancestors in common dating back approximately 600 years ago). —is thus critical to a GWAS outcome. Prioritizing ”cohort matching” in a genetic study ensures that the test sample closely aligns with the reference cohort in key factors such as ancestry, age, geography, and social or environmental influences. This alignment minimizes confounding and enables a more accurate interpretation of genetic effects on traits. [^6] For example, (Mostafavi et al., 2020) find that genetic data prediction accuracy varies within an ancestrally-homogeneous white British cohort, e.g., weights from a BMI GWAS conducted in a younger cohort have higher prediction accuracy via a PGS in younger individuals than older individuals.
3. Background: Current Protections
Figure 1. Four Pillars of genetic data collection and their regulatory considerations. Data collectors within each pillar operate with different collection strategies and intent to store, use, and transfer the data. However, there is only a patchwork of federal and state-level legislation that governs each Pillar. The resulting regulatory gaps lead to leaky protections through which genetic data—or knowledge derived from genetic data—can be used in other pillars with little to no oversight. Real-world examples [^8] Orchid Health is a private company that uses publicly available GWAS weights to compute polygenic risk scores (PGS) of embryos. Nucleus Genomics is also a company that similarly uses whole genome sequencing to compute PGS to predict health and behavioral traits. of leakage are indicated by red arrows.
In this section, we outline the foundation of today’s genetic data collectors and the current regulatory and legislative environments that they are subject to. To better understand the vast landscape over which genetic data is being collected, we build upon the Four Pillars (Wan et al., 2022) of genetic data collectors. We categorize these data collectors by their motivations for collecting genetic data and their relationships with the individuals from whom the data is collected. We define the updated Four Pillars as the following:
  • Forensics, Tracking, Surveillance: Entities focused on monitoring, identifying, and analyzing genetic data to support public health outcomes, security, and law enforcement objectives.
  • Direct to Consumer (DTC): Entities developing products to provide consumers with direct access to genetic insights.
  • Clinical: Entities requiring medical professionals’ oversight to conduct tests, often enabling insurance coverage and generating clinically actionable insights.
  • Research: Entities engaging in research and development (R&D) and basic research across private, public, and academic sectors. This includes public consortia, biobanks, and collaborations with DTC entities.
The “genetic data ecosystem” encompasses not only the entities within the Four Pillars but also the broader network of legislatures, regulators, and individuals involved in or affected by genetic data collection and use. By outlining this ecosystem, we aim to shed light on the regulatory gaps that enable genetic data leakage across pillars. We provide real-world examples of this leakage in Figure 8.
3.1. Federal protections
3.1.1. Genetic Information Non-Discrimination Act (GINA)
GINA was enacted in 2008 and prohibits genetic discrimination by employers or health insurers. Specifically, GINA bars employers from making hiring, firing, or promotion decisions based on genetic information (and information disclosed about relatives) and bans health insurance providers from denying coverage or charging higher premiums based on genetic information including genetic predisposition for disease (Congress, 2008). GINA does not apply to life, disability, or long-term care insurance (Green et al., 2020) or any other venues outside of employment and health insurance (e.g., education). The US Equal Employment Opportunity Commission (EEOC) regulates workplace discrimination for employees and thus, is responsible for enforcing GINA in the workplace. Enforcement for health insurers falls under the purview of several federal agencies, including the Department of Health and Human Services (HHS); Department of Labor; Centers for Medicare and Medicaid Services; and the Department of the Treasury (Protections, OHRP).
3.1.2. Health Information Portability and Accountability Act (HIPAA)
HIPAA, signed into law in 1996, grants security and privacy protections for patient health information (PHI). HIPAA applies to covered entities that maintain this information, such as healthcare and health insurance companies, but does not apply to many other entities that might retain health information, including search engines, medical information sites, or dating sites (Citron, 2020). Importantly, HIPAA does not apply to two of the largest DTC genetic testing companies 23andMe and AncestryDNA (Sklar, 2020). HIPAA does not cover deidentified data: PHI data with unique identifying information (names, address, phone numbers, dates, etc.) removed do not fall under HIPAA. Research consortia, biobanks, and research collaborations with DTC and clinical entities primarily use genetic data that is considered deidentified, and so HIPAA generally does not apply. However, the relational aspect of genetic data has motivated calls to “consider genomic data as, in principle, always identifiable” (Bonomi et al., 2020). The inability to completely strip data of all identifiable features introduces critical vulnerabilities for genome owners and exacerbates the risk of data leakage across the Four Pillars.
3.1.3. Food and Drug Administration (FDA)
The FDA “is responsible for protecting the public health by ensuring the safety, efficacy, and security of human and veterinary drugs, biological products, and medical devices” (Commissioner, 2023). In the context of DTC genetic testing companies, the FDA only regulates well-defined high risk medical tests with clinical actions, such as 23andMe’s health tests for diseases like Parkinson’s, breast cancer, and late-onset Alzheimer’s. If the FDA approves a specific test, Centers for Medicare and Medicaid Services then choose to approve the test for insurance coverage (Daval and Kesselheim, 2023). Most genetic tests are not covered by Medicare, particularly DTC tests.
3.1.4. Federal Trade Commission (FTC)
The FTC enforces federal laws that protect consumers from “fraud, deception and unfair business practices” (Commission, 2025). The FTC has warned DTC genetic testing companies that the results they return to customers must be backed by “reliable science”, and also warns companies against making exaggerated claims about the use of AI in their products (Jillson, 2024). Additionally, the FTC remains vigilant of deceptive practices regarding privacy policy changes and “dark patterns” designed to coerce consumers into consenting to data sharing [^9] In 2023, the Commission fined 1Health.io/Vitagene \$75,000, claiming they “left sensitive genetic and health data unsecured, deceived consumers about their ability to get their data deletes, and changed its privacy policy retroactively without adequately notifying and obtaining consent” (Commission, 2023).. The power and broad scope of the FTC—and their stated intentions of cracking down on DTC genetic testing companies (Jillson, 2024)—make it a major player in the regulatory space, and it will likely grow in importance if the federal government prioritizes genetic data privacy.
3.1.5. The Common Rule
This is a federal policy that protects human subjects in research settings (Sciences et al., 2014). It was first codified in 1981 by the Department of Health, Education, and Welfare (now HHS) but adopted more widely by other federal agencies in 1991. The Common Rule institutionalizes protections for vulnerable participants, protections overseen by Institutional Review Boards (IRBs); IRBs are groups “formally designated to review and monitor biomedical research involving human subjects” (Research, 2024). Revisions enacted in 2017 particularly focus on informed consent for participants, but do not cover deidentified data (Menikoff et al., 2017; Protections, OHRP). IRBs ensure that consistent protection of human subjects are maintained and are necessary for research to begin and proceed if involving human participants. The maintenance of the Common Rule and HIPAA, in the case of medical data, is performed by IRBs.
3.2. State protections
Figure 2. Beyond HIPAA and GINA, state-level genetic discrimination and privacy legislation varies widely in the United States. Left: States with genetic anti-discrimination legislation, distinguishing between states with both health insurance and employment protections (red) and those with health insurance protections only (orange). Right: states with general privacy laws that include genetic data (blue) and genetic privacy-specific laws (green).
State laws offer far more protections than their federal counterparts, but a patchwork of these laws means a patchwork in protections, as shown in Figure 2. Here, we focus on two categories of laws: genetic anti-discrimination laws and data privacy laws.
3.2.1. Genetic Anti-Discrimination Laws
Most states have passed some version of a genetic anti-discrimination law (initially prompted by the passage of GINA) although the sector, scope, and strength of the laws differs between states. Some states offer fewer protections than GINA: Georgia, for example, has a statute outlawing genetic discrimination by health insurers, but not employers. Other states offer more protections than GINA. California’s CalGINA (Legislature, 2011) outlaws genetic discrimination in several sectors: employment, health insurance, housing, and by any state agency or entity receiving state funding (such as emergency medical services). Florida includes life, long-term care, and disability insurers (of Representatives, 2020). Many states (including Massachusetts and New Mexico) that include life insurance in their non-discrimination laws exempt companies if genetic data reliably, and based on “sound actuarial principles” (Legislature, 2025), gives information relating to mortality or morbidity (see Section 5.4 for a detailed case study).
3.2.2. Genetic Data Privacy Laws
Many states have genetic data privacy laws, either as a standalone law or as part of a broader data privacy law. These laws [^10] Some of these laws have also led to a change in company services in that state. For example, the passing of Illinois’ BIPA (Biometric Information Protection Act) resulted in Amazon and Google removing certain services of Alexa and Google Photos respectively after lawsuits which claimed that the company violated BIPA. were generally passed after GINA with the rise of DTC genetic testing companies. Standalone genetic data privacy laws tend to be written specifically to cover DTC genetic testing companies and are often extensions of genetic non-discrimination laws passed around the time of GINA.
Both data privacy and genetic privacy laws have similar provisions (Institute, 2025), including a customer right to request deletion of genetic data and samples and the necessity of consent for data transfers. Some states allow consumers to bring lawsuits for violations of the privacy law (a private right of action). In general though, privacy laws do not apply to deidentified genetic data (as per HIPAA). An important contrast to the US federal and state bills on privacy is the EU General Data Protection Regulation (GDPR). GDPR not only regulates all personal data within the EU but also the transfer of data outside of the EU, data minimization, and compliance. Genetic data is labeled as “special category data” in GDPR as it has many identifiers associated with it.
4. Risk Assessment Framework
Our goal in this section is to develop a risk assessment framework for genetic data governance. Similar to the aims of the Blueprint for the AI Bill of Rights (Park, 2023), we motivate our framework with three central questions: (1) What values (i.e., moral principles and civil liberties) should be preserved? (2) What are the vulnerabilities in the current system that can compromise these values? (3) What are the specific harms that can result from the vulnerabilities?
4.1. What values should be preserved?
We seek to identify values that are future-proof and can be encoded in a system of genetic data collection. By examining the Four Pillars, six key values emerge:
  1. (1)
    Right to action: The individual, and the individual only, has the choice to submit (and the freedom to not submit) their genome [^11] We use “genome” to refer to a physical DNA samples of an individual and any sequencing/genotyping data..
  2. (2)
    Ownership of the genome: The individual owns their genome and therefore controls the usage of their genome, including who has access to it and for what reason.
  3. (3)
    Right to privacy: The individual has a right to privacy with regards to their genome and inferences made from their genome.
  4. (4)
    Right to knowledge: The individual has a right to know or not know about inferences made from their genome.
  5. (5)
    Protecting opportunities for advancement: Genetic data should not be used to deprive the individual of opportunities in any domain, including (but not limited to) education, access to financial tools, health insurance, housing, social services, and reproductive choices.
  6. (6)
    Benefits of inclusion: The Belmont principle states “those who bear the burdens of research (i.e., those who are exposed to the discomforts, inconveniences, and risks) should receive the benefits in equal measure to the burdens”(U.S. Department of Health & Human Services (1979), HHS). We believe that this applies to individuals and their genetic data. [^12] Variant Bio is a biotech company that collects genetic data from Indigenous groups from around the world and participates in revenue sharing with the communities they collect data from.
4.2. What are the vulnerabilities in the current system that can compromise these values?
The current genetic data ecosystem has many vulnerabilities that compromise the values listed above. These include:
  1. (1)
    Unsettled science: The role of genetics in shaping complex traits is poorly understood. As a consequence, the deterministic nature of genetics may be overstated, particularly for behavioral and cognitive traits.
  2. (2)
    Rapid evolution of genetic data/methods: As genetic data collection, sequencing, and methodologies quickly evolve, legal protections fall behind. For example, education discrimination is not included in GINA, which was passed five years before the first large GWAS on educational attainment (Rietveld et al., 2013).
  3. (3)
    Guilt by association: DNA databases used for criminal investigations can impact not only the individuals whose DNA is stored, but also their biological relatives, as genetic information is shared among family members.
  4. (4)
    Geographical legislative patchwork: Genetic anti-discrimination and privacy legislation widely differs between states. Legal policy in each state for data removal, third-party sharing, and many other aspects of genetic data are unclear. Individuals who move between states or share data with entities in other states can be affected by this dependency on state specific policy. Aside from patchwork policy, many state policies do not sufficiently protect genetic data specifically.
4.3. What are the harms that arise as a result?
Lastly, we will outline several consequences that could arise from violations of our key values through the above vulnerabilities. These downstream effects, which we refer to as harms, can affect an individual and their immediate genetic relatives and can also have a broader effect on communities and the population at large.
  1. (1)
    Leakage to the family: The genetic dragnet means that any conclusion about an individual’s genetic trait can be linked back to their relatives, even if no action is taken by their relatives in either (1) submitting data or (2) opting in to receive information. Any derived secondhand knowledge can immediately impact insurance (health, life, etc.) and medical treatment plans, as well as compromise identity.
  2. (2)
    Loss of anonymity: It is straightforward to identify some traits about an individual from their genetic data, and subsequently from their genetic relatives. Common examples include: race, gender, ethnicity, and markers for certain diseases. Leaky genetic data interfaces can thus compromise information about an individual, their genetic relatives, and unborn children that the individual (and implicated relatives) may wish to maintain as private. As the refinement of reference datasets continues and data collection grows [^13] It has been estimated that, in a sample of three million European-descent Americans, 99% of individuals would have a least one third cousin in the dataset (Erlich et al., 2018)., anonymity becomes even more elusive.
  3. (3)
    Loss of Data Control: Private institutions offering genetic tests control the entire lifecycle of individuals’ data from collection and storage to monetization. This control creates significant transparency issues, as it is often unclear how data is monetized, which parties benefit, or what agreements include an individual’s data. Users may opt out of participation, but unregulated data flows ensure that their genetic footprint persists indefinitely. This lack of transparency is particularly concerning in cases of bankruptcy or acquisition. For example, in 2024, Tempus AI acquired Ambry Genetics, a company that sequences genetic data for 400,000 patients annually (Tempus, 2025). Tempus stated its intent to ”leverage this data and augment Tempus’ current data offering,” raising questions about how patient data is repurposed and monetized in such deals. Ultimately, an individual’s decision to submit genetic data for testing often entails relinquishing full ownership of their genetic information. Without robust safeguards, privacy risks emerge from opaque data practices and leaky interfaces, leaving individuals vulnerable to misuse or unauthorized access to their genetic data.
  4. (4)
    Misinformed Actions: The interpretation of genetic test results and the subsequent actions taken by individuals are profoundly personal, often influencing lifestyle changes, significant financial decisions, and critical medical choices. For instance, individuals with BRCA1 gene mutations indicating an elevated breast cancer risk may opt for preventive interventions. These decisions are particularly sensitive as the interpretation of genetic data evolves alongside advancements in scientific methods. However, the relational nature of genetic data complicates individual autonomy and access to information. It is possible for someone to receive conclusions about themselves indirectly through the test results of genetic relatives. This dynamic, combined with the widespread availability of unregulated genetic tests and the evolving nature of genetic science, increases the potential for harm if individuals act on information that is incomplete, inaccurate, or probabilistic rather than deterministic.
  5. (5)
    Financial impact: Individuals may face financial repercussions if insurance or legal policies fail to provide adequate protection or if coverage is denied. Beyond medical procedures, financial impacts can also arise on a case-by-case basis, such as in situations where genetic data is held for ransom, legal defense is required, or individuals are forced to purchase direct-to-consumer tests (e.g., in order to purchase a life insurance policy).
5. Case Studies
We apply our risk assessment framework to four case studies which highlight regulatory gaps across the Four Pillars. These case studies, ranging from past events to speculative scenarios, all involve privacy loss or potential discrimination. We choose case studies that illustrate genetic data leakage in both, a pillar-to-pillar and one-to-many pillar settings.
5.1. Genetics and Education
Polygenic scores (PGS) are popular in the social/behavioral sciences for predicting social outcome traits, such as educational attainment (EA; number of schooling years completed by an adult). We will use EA as the example throughout this case study, but note that there is also interest in predicting standardized testing scores, performance in mathematics, and other traits with substantial environmental influences.
A common metric for assessing PGS accuracy is the percentage of trait variance it explains: higher percentages indicate better predictive performance. A recent 23andMe EA PGS, based on data from over 3 million customers of European descent, explains 12-16% of the variance  [^14] As a useful comparison, mother’s education explains 15% of the variance in EA (Lee et al., 2018) in educational attainment (EA) (Okbay et al., 2022). This effectively means that 50-70% of individuals with PGS scores in the top 10% for EA are predicted to graduate college. However, the PGS accuracy significantly decreases when applied to African American customers. This is an example of the commonly observed “portability problem” (Martin et al., 2017), where a PGS derived from GWAS in one population predicts poorly in another due to confounding [^15] Confounding in GWAS can be genetic (non-causal variants correlated with causal ones) or environmental (non-causal variants correlated with causal environmental factors), leading to potential statistical artifacts in effect sizes..
There have been several calls to use EA PGS to inform education policy. Harden et al. propose the use of math-performance PGS to identify “leaks” in the education system: for example, by identifying high math PGS students who perform poorly, they claim educators could pinpoint why and how students are failing to reach their potential [^16] An example they use: 31% of high PGS students in good schools take calculus, compared to 24% of the same-scoring students in poor-performing schools. (Harden et al., 2020). Plomin & von Stumm take it further: they use the term “precision education” (akin to precision medicine) to propose a tailor-made, individualized education that is genetics-informed (R and S, 2018). Statements like this, combined with statements such as “students with higher polygenic scores for years of education have, on average, higher cognitive ability, better grades and come from families with higher SES [socioeconomic status]” (Smith-Woolley et al., 2018) are cause for concern because they invoke a sense of genetic determinism. However, other predictors (parents’ educational status, socioeconomic status) explain similar amounts of variance in EA (Lee et al., 2018; Morris et al., 2020) and—unlike DNA—are mutable through social policy changes.
Several of our values would be violated if children were required to submit their DNA (Right to Action) or educational opportunities were denied to children based on their genetic potential (Opportunities for Advancement). Through the vulnerabilities of unsettled science and the rapid evolution of genetic methods, harms such as leakage to the family can occur and affect not only children but their families and future.
5.2. DTC Genetic Data Brokerage and Leakage
Several FDA-approved 23andMe tests17171723andMe tests with FDA 510(k) clearance are verified as substantially equivalent to existing devices or methods. Only health-related tests, such as those identifying genetic risks, carrier status, or drug responses, qualify for this review. provide customers with results containing sensitive information, including carrier screening reports, genetic health risk assessments, and BRCA1/BRCA2 variant analyses. In addition, they offer genealogy (to identify relatives) and ancestry services. In an attempt to protect customer privacy, 23andMe offers an opt-in/opt-out policy when it comes to sharing their samples in studies with third-parties.
In 2023, hackers accessed 14,000+ 23andMe user accounts by using recycled login information (credential stuffing). Due to the relatives linkage feature in genetic data reports, the hackers were able to then access information about an additional 5.5M users (23a, 2023b); there were also reports that the hacker targeted Ashkenazi Jewish and Chinese ancestry profiles (Carballo et al., 2024). Among the many values violated, the Right of Privacy and Ownership of the Genome are particularly relevant to this case as 23andMe did not immediately notify compromised users or the public. Several class action lawsuits were filed (e.g., (for the Northern District of California, 2023)), with plaintiffs complaining that 23andMe failed to adequately protect their sensitive information. 23andMe responded by enforcing multifactor authentication (MFA) of user accounts, which was previously optional (Whittaker, 2023), and a serious technological vulnerability. A key gap that allowed for this issue to occur is the lack of consensus around security responsibility, which resulted in the disproportionate and harmful targeting of underrepresented groups. The 23andMe breach highlights the consequences of the relational nature of genetic data, where an individual’s information can be compromised through poor public protections, even if they personally take precautions. As one Reddit user quipped “Your genetic data is only as secure as your relatives’ passwords.”
5.3. Genetic Data Collection of Detained Noncitizens
A recent Georgetown University report from the Center of Privacy and Technology (CPT) examines the U.S. federal government’s practice of collecting DNA from detained migrants (Glaberson et al., 2024). This practice, which began with the 2005 DNA Fingerprint Act and expanded significantly in 2020—a mandate that the Department of Homeland Security must collect DNA from all detainees, even briefly detained [^18] States have varying rules about DNA collection from suspects and detainees: 28 states allow DNA to be taken from someone before they are convicted of any crime and most states restrict DNA collection to cases involving felony charges or convictions, rather than allowing collection for any arrest or detention (Samuels et al., 2012).—has led to a dramatic increase in detainee representation in CODIS (the federal DNA database), rising from 0.21% in 2019 to 9.21% in 2023. By 2020, approximately 25,000 noncitizens were added to the database under the “detainee” classification.
The CPT report highlights several concerning aspects of this practice, particularly its disproportionate impact on migrants of color and issues of consent. In our Risk Assessment Framework, this immediately violates the Right to Action and Right to Knowledge. Many migrants undergo DNA collection without understanding its implications, sometimes believing it to be a COVID-19 test or submitting under threat of criminal prosecution. The report argues that this practice violates the Fourth Amendment in collecting DNA from detainees without probable cause. Additionally, even though CODIS was designed to be privacy-protective by collecting only 20 markers—a tiny slice of the genome believed to be medically neutral—this limited data can still identify relatives through partial matching and, as an analysis 25 years later would show, are slightly informative of disease risk (Bañuelos et al., 2022). The expansive collection of genetic data has far-reaching implications through the guilt-by-association vulnerability. Since CODIS records are difficult to expunge, they can restrict Opportunities for Advancement not only for detainees but also for their relatives and descendants.
5.4. Underwriting Life Insurance with AI and Genetic Data
Life insurance underwriting is the process whereby an insurance company uses personal and health information to assess the risk of insuring an applicant. The relationship between applicants and insurance companies is already fraught: (Devnos et al., 2016) find that patients are more likely to share their genetic info with Google than insurance companies. The future of life insurance underwriting is expected to become increasingly computational and automated via the usage of AI and the collection of more personalized, individual genetic data (Filabi and Duffy, 2021; Balasubramanian et al., 2021; Rothstein, 2018; Koleva-Kolarova et al., 2022). While medical records and demographic factors can be used for mortality analysis, the inclusion of genetic factors can mean that risk prediction can be performed much earlier in an applicant’s life without necessarily the same amount of records as an older individual, emphasizing the risk of genetic discrimination (Karlsson Linnér and Koellinger, 2022) based on the potential to have risk factors (e.g., a genetic risk for high blood pressure versus clinically-measured high blood pressure). Life insurance is not covered by GINA and life insurance companies can access medical records (which may include genetic test results) as part of an application. Bills have been introduced in several states that would restrict the use of genetic information in underwriting, but these efforts largely failed: As of 2022, out of thirty-seven proposed bills across all states, three were introduced, eight were signed by the governor, demonstrating the geographical dependency of policy-based protection (Rickard, 2022).
Life insurance companies may be interested in using polygenic scores (PGS) to assess an individual’s risk for various diseases. Indeed, in early 2024, U.S.-based life insurance company MassMutual and U.K.-based Genomics plc announced a partnership, offering free genetic testing to MassMutual’s life insurance customers (MassMutual, 2024). However, the press release stressed that MassMutual would not receive individual results and that current premiums/policies would be unaffected (MassMutual, 2024). This example particularly highlights the leaky interface between pillars, where the Right of Privacy and Right of Action are called into question. AI insurance underwriting algorithms already suffer from racial biases (Lee, 2022). That PGS suffer from the portability problem (itself caused by systematic biases in training datasets) means that these biases could be perpetuated in genetics-informed life insurance underwriting. Additionally, given the relational aspect of genetic data, genetic-based underwriting could affect other biological relatives’ applications, harming not only relatives but any others with the same genetic markers that are associated with mortality risks. The potential harms of using genetic data as a part of underwriting in life insurance are one such example of violations of genetic privacy with the rapid evolution and usages of genetic data with unsettled science.
6. Recommendations
We argue that for public infrastructure and private companies to productively and ethically make use of genetic data, several amendments need to be made to the current genetic data governance system. Our three recommendations address open privacy concerns, legislative scoping for policy changes, and best practices for bodies handling genetic data.
6.1. Recommendation 1: Redefining Genetic Data
Issue: Legal policy surrounding genetic privacy notably excludes deidentified or anonymized data from protection.
Recommendation: Given that we argue that genetic data is unique compared to any other identifying data, we suggest genetic data be defined using the following language:
“Genetic data refers to any information relating to an individual’s genetic characteristics, including but not limited to DNA or RNA sequences, gene expression profiles, or any data derived from a biological sample—including that of a relative—regardless of format. This data is always considered identifiable by its nature—or personally identifiable information (PII)—as it pertains to unique biological attributes that can potentially be linked to a specific individual, their biological relatives, or identifiable group. For the purposes of this definition, any de-identified, pseudonymized, or anonymized genetic information shall be treated as genetic data, regardless of measures taken to mask individual identities, recognizing that the inherent characteristics of genetic information may enable re-identification through advanced technological or data cross-referencing methods.”
6.2. Recommendation 2: Extending Protections for Genetic Discrimination
Issue: GINA is a vital piece of federal legislation that protects against genetic discrimination in employment and health insurance domains. However, other domains in which there exists potential for genetic discrimination are not protected by GINA: other insurance domains (life, long-term care, disability), housing, and education.
Recommendation: We recommend that additional federal laws should be enacted to extend GINA’s coverage beyond employment and health insurance. The most comprehensive state genetic anti-discrimination law is California’s CalGINA, which includes housing, mortgage brokerage, education, and more. However, despite these extensions, CalGINA does not cover life, disability, or long-term care insurance. We suggest that extension bills for GINA should explicitly cover life, disability, long-term care, as well as education and any other opportunities for advancement to be protected against genetic discrimination. It should be written in such a way that prohibits any barrier to opportunity, even those not anticipated at the time of writing. Additionally, legislation should be written to explicitly bar genetic risk for complex traits known to have significant environmental influences (such as cardiovascular disease) from being considered a preexisting condition.
6.3. Recommendation 3: A Genetic Data Regulation Framework
Issue: Current regulations were designed to govern one application of genetic data—or Pillar—at a time. This has led to “leaky protections”, where the use of genetic data in one Pillar can affect opportunities and decisions in other Pillars (e.g., clinical tests being used in life insurance).
Recommendation: To eliminate leaky protections, we suggest a comprehensive and uniform regulatory framework that encompasses all usage domains whether genetic data is collected, analyzed, or used in any type of inference. Below, we build upon specific values inherent in privacy rights to ensure that responsibility is placed on the organizations that hold and analyze data, rather than the individuals from whom it was collected (Solove, 2023).
  1. (1)
    Entity approval: Entities—whether they be universities, health care providers, hospital systems, or corporations—that collect or house any type of genetic data (DNA, RNA, or even genetic test results) must have prior regulatory approval. Approval requires a clear commitment to the basic rights of the individuals whose genetic data the entity will collect or own, such as data privacy, security, the right to deletion, etc.
  2. (2)
    Test approval: After entity approval, we propose regulations for the inferential tests, specifically those that return results to either an individual or a medical practitioner for any actionable result. At a minimum, entities would be required to publicly release detailed white papers for each test that detail laboratory procedures, quality control steps, inferential models used, presentation of results, and any other necessary details that would allow one to recreate the analysis if given the same data.
  3. (3)
    Powers given to the individual
    • Individuals should be able to request data removal. This would include specifically (1) the destruction of the physical sample, (2) deletion of data, (3) removal of any identifiers, and (4) that any removed data no longer influence the results of any downstream models.
    • Third-party data transfers should only occur between companies authorized to collect or own genetic data, with recipient companies subject to the same regulations. We recommend providing individuals with the option of blanket opt-out from data transfers and specific opt-in choices for each third party. Individuals must be notified of the transfer and given a reasonable time to opt-out.
    • Research usage is common among DTC companies, academia, and hospitals, where data provided per a test can be pooled together as a dataset for research (both internally and via external collaborations). We suggest an opt-out strategy here for individuals, with a specification on whether they consent to their data becoming publicly available in any form (e.g. public datasets, open sourced GWAS weights, trained models, etc).
    • Incidental genetic discoveries can be possible in research use cases where data for a particular test was used for other tests. In these cases, where knowledge of a particular trait can also be harmful and uninformative to the individual, we suggest an opt-in strategy for an individual to blanket choose if they wish to hear any secondary discoveries, or forward any secondary discoveries to their medical practitioner, to be enacted on at their discretion.
  4. (4)
    Bankruptcy and Acquisition: As detailed by third-party data transfers, the owning of genetic data would also apply to companies who acquire any genetic data as part of assets through acquisitions or bankruptcy. Thus, all entities involved must already have approval to handle genetic data. In the rare event that there are no entities to handle the data, we suggest a similar protocol structure as nuclear waste (Doctorow, 2008), where all physical sample, data, and models are subsequently destroyed and cannot be recovered.
7. Conclusions and Open Challenges
In this paper we outline the genetic data ecosystem and today’s state of public protections. We identify critical gaps in current regulatory federal and state-level frameworks that prevent the productive, ethical, and safe adoption of genetic data for broad societal applications. To address this, we propose a novel risk assessment framework that offers key public protections and preserve’s individuals personal liberties, and finally make three concrete policy recommendations.
While our policy recommendations focus on enhancing user protections, we also identify three key challenges in regulating genetic data usage that fall outside the scope of our work and that would benefit from further research.
7.1. The relational theory of privacy.
7.1.1. Should genetic data submissions or inclusion into genetic datasets require consent from biological relatives?
When an individual does a genetic test they are, indirectly, testing (part of) their relative’s genome. What (if any) consent or privacy measures should be taken to protect their biological relatives who may not consent to such a test? Additionally, what should happen upon death? Do children “own” their parents’ genome (and what ownership would, say, a full sibling have?) and can they make decisions (e.g., to delete the data)?
7.1.2. Who has ownership of children’s DNA?
Do parents have a “fiduciary duty” to protect their child’s genetic data and protect them from the system vulnerabilities that we have outlined? Part of this fiduciary duty would involve whether a genetic test should be taken in the first place, e.g., the difference between a medically-necessary test and a DTC genetic test19191923andMe’s policy for minors is as follows: individuals under 18 that give assent and have a parent/guardian’s permission can submit their sample (23andMe, 2024a). for some insights into the child’s ancestry and genealogical relatives. The relational aspect of genetic data also comes into play here; for example, should consent be required from both parents to sequence their child’s genome? See (Bala, 2023) for a thorough discussion of child’s genome ownership.
7.2. International genetic data transfers.
Our paper focuses on American regulatory and legislative bodies and American companies. In the DTC genetic testing space, individuals may be less likely to trust foreign companies: Nucleus Genomics [^20] https://mynucleus.com proudly states on their home page that customer genomes are sequenced in the U.S. and the data is never sent/stored overseas21212123andMe also genotypes samples in the U.S. (23andMe, 2024b), but this is not a selling point of theirs.. Should there be restrictions on genetic data transfer between countries? For example, one hypothetical scenario might involve a political refugee who has medically-necessary genetic testing done within the U.S, but because of leaky data transfer rules, genetic information is transferred to their home country where it might be used to identify family members now at risk for persecution.
7.3. The deployment of AI models for genetic data.
Our paper has largely shied away from discussions about genetic inference and AI for two reasons: (1) deployment of AI technologies for genetic inference is nascent and, most importantly, (2) risks and vulnerabilities exist even with current simple linear models for genetic inference. AI will, we believe, exacerbate the risks we have outlined as they will likely result in seemingly better trait prediction by modeling nonlinear interactions and taking advantage of gene-environment correlations(Yang et al., 2024). However, the uncertainties associated with the basic scientific inference (that we outline in Section 2) will be compounded because of the lack of interpretability and complexity of AI models.